OpenStudy (anonymous):

Evaluate the integral e^-0.01x-6 dx from 0 to 100

zepdrix (zepdrix):

Is the -6 in the exponent?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large \int\limits_0^{100}e^{(-0.01x-6)}dx\] Using laws of exponents, we can rewrite our exponential like so,\[\large e^{(-0.01x-6)} \qquad = \qquad e^{-0.01x}\cdot e^{-6}\] Notice that e^{-6} is a constant with respect to x, so we can simply pull it out of the integral, giving us,\[\large e^{-6}\int\limits_0^{100}e^{-0.01x}dx\] Understand what I did so far?

OpenStudy (anonymous):

yes I do

zepdrix (zepdrix):

Do you understand how to evaluate this integral? \[\large \int\limits e^{2x}dx\]

OpenStudy (anonymous):

I think so. Is the answer -100/e^7 + 100/e^6 ?

zepdrix (zepdrix):

Yes, very good! :)

OpenStudy (anonymous):

thank you!